# moonboot: An R Package Implementing m-out-of-n Bootstrap Methods

### v2

@@ 1   Introduction
The bootstrap is a resampling method introduced by Efron (1979), which repeatedly simulates samples of the same size $n$ as the observed data by $n$-fold drawing with replacement. Let us call this method n-out-of-n bootstrap, in order to distinguish it from different sampling schemes. For each of these samples, the estimator $T$ of interest is computed, thereby simulating a sample $T_{1}^{*},\ldots,T_{R}^{*}$ for the distribution of $T$, where $R$ is the number of bootstrap repetitions. For sufficiently smooth estimators $T$, this bootstrap distribution asymptotically approaches the true distribution of $T$ (Giné and Zinn, 1990; Shao and Tu, 1995), and it can thus be used to construct confidence intervals for $T$. The coverage probability of confidence intervals based on flipping the quantiles of the bootstrapped distribution $T_{1}^{*},\ldots,T_{R}^{*}$ at the point estimate (“basic bootstrap”) can be shown to be first order accurate, i.e. up to $o(n^{-1/2})$. If bias correction and acceleration methods are applied (“$BC_{a}$ bootstrap”), they are even second order accurate, i.e. up to $o(n^{-1})$ (DiCiccio and Efron, 1996; Davison and Hinkley, 1997). Compared to other methods for estimating confidence intervals, the bootstrap is more versatile because it neither relies on a likelihood function, nor does it require asymptotic normality. The bootstrap has thus become a standard technique for estimating confidence intervals, and it is even included in vanilla R with the package boot (Canty and Ripley, 2021), which implements five different methods in boot.ci().
@@
There are, however, bootstrap inconsistent estimators, i.e., estimators for which the n-out-of-n bootstrap fails to yield confidence intervals with an asymptotically correct coverage probability. Examples include extreme order statistics (Bickel et al., 1997), the Grenander estimator of a monotonous density (Sen et al., 2010), Chernoff’s estimator of the mode and Tukey’s shorth (Léger and MacGibbon, 2006), or Chatterjee’s rank correlation index (Lin and Han, 2024; Dalitz et al., 2024). See the discussion by Lin and Han (2024) for a literature review and further examples. A possible workaround in these cases is the m-out-of-n bootstrap, which samples only $m<n$ observations. This can be done with or without replacement, which Bickel et al. (1997) called the “m-over-n ($m/n$) bootstrap” and the “n-choose-m bootstrap”, respectively. Politis and Romano (1994) have shown that the n-choose-m bootstrap works under much weaker conditions than the n-out-of-n bootstrap, provided $m$ is chosen such that, for $n\to\infty$, $m\to\infty$ and $m/n\to 0$. The only condition is that the estimator, suitably scaled by some factor $\tau_{n}$, possesses a limit distribution, whereas no smoothness of the estimator or uniformity of the convergence is required. If the sampling is instead done with replacement, Bickel et al. (1997) have shown that a similar result requires additional restrictions on the estimator. Sampling without replacement thus works under weaker conditions than sampling with replacement, and, in the present article, we therefore primarily consider sampling without replacement, unless otherwise stated.
@@
The wider applicability of the m-out-of-n bootstrap comes at a price, though. One drawback is that the scaling factor $\tau_{n}$ must be known. Bertail et al. (1999) explained that it must be chosen such that $\tau_{n}^{2}\operatorname{Var}(T)$ converges to some constant, and they also suggested a method to estimate the scaling factor with another bootstrap. For root-$n$ consistent estimators, it is $\tau_{n}=\sqrt{n}$, but not all estimators are root-$n$ consistent, especially if they are bootstrap inconsistent, and even if root-$n$ consistency is conjectured it can be difficult to prove (Lin and Han, 2024). Moreover, the m-out-of-n bootstrap can have less than first order accuracy: Politis and Romano (1994) gave a theoretical example with accuracy $O(n^{-1/3})$ for an optimal choice of $m$, and even worse for other choices of $m$. That the accuracy depends on the choice of $m$ was demonstrated, too, in a simulation study by Kleiner et al. (2014). And for the sample quantiles and the m-over-n bootstrap, Arcones (2003) has shown that the choice $m\propto n^{2/3}$ is optimal. This raises the problem how to choose $m$, and several heuristics have been suggested in the literature (Politis et al., 1999; Götze and Račkauskas, 2001; Bickel and Sakov, 2008; Chung and Lee, 2001; Lee and Yang, 2020).
@@
Due to these shortcomings, the usual n-out-of-n bootstrap is generally preferable for bootstrap consistent estimators. However, there are bootstrap inconsistent estimators and there is thus need for an R package that facilitates the application of the m-out-of-n bootstrap by providing the required functions. To this end, we present a new package moonboot, evaluate the algorithms implemented therein by means of Monte Carlo simulations, and give recommendations and use cases for their application.
@@ 2   The m-out-of-n bootstrap
Let $X_{1},\ldots,X_{n}$ be i.i.d. random variables, $\theta$ be some parameter of their underlying distribution, and $T_{n}=T_{n}(X_{1},\ldots,X_{n})$ be an estimator for $\theta$. Let us additionally assume [^1] Politis and Romano (1994) did not make this assumption, but it is necessary for the estimation of the scaling factor $\tau_{n}$ according to Bertail et al. (1999) that $E(T_{n}^{2})<\infty$. The m-out-of-n bootstrap requires that, for some scaling factor $\tau_{n}$ which behaves for $n\to\infty$ as
@@
Note that condition (1) implies that $\tau_{n}\to\infty$, and thus the convergence (2) requires that $T_{n}$ converges to $\theta$ in probability, because it follows from (2) that
[…]
@@
Sherman and Carlstein (2004) observed that the scaling with $\tau_{m}/\tau_{n}$ in (4) can be omitted if the interval is centered around one of the $T_{m}$ instead of $T_{n}$. They suggested to use the first $m$ samples for computing $T_{m}$, but this is an arbitrary choice and thus even the location of the confidence interval is not uniquely determined by the data. Moreover, the lack of scaling increases the interval length. The greater length is compensated by the higher voltility of the location so that the coverage probability still is close to the nominal value. This does not hold, however, for drawing without replacement as $m$ approaches $n$, and a compromise must be made between small interval length and approximate coverage probability. Sherman and Carlstein (2004) suggested a heuristic method for choosing $m$ on basis of a double bootstrap, and they gave the rule of thumb to choose $m$ between $n^{1/2}$ and $n^{2/3}$. However, due to the greater interval length, this method should only be used as a last resort if the convergence rate $\tau_{n}$ is unknown and cannot be estimated from the data. In moonboot, this method is implemented by mboot.ci(..., method="sherman") and the method for estimating $m$ by estimate.m.sherman().
@@ 2.1   Estimation of $\tau_{n}$
and $\beta$ can be estimated with a least-squares fit. For the choice of the test values $m$, Bertail et al. (1999) sugested $m_{i}=n^{\beta_{i}}$, but wrote that “the difficult problem of choosing the $\beta_{i}$’s requires more work on a case-by-case basis”. This casts doubt on the whole method, because removing the necessity of a detailed asymptotic analysis of the estimator is the whole point of estimating $\tau_{n}$ from the data. Indeed, we have not found a sequence of values for $\beta$ that worked equally well in all of our examples. We thus leave the choice of the $\beta$-sequence as an option to the user and use the default of five values between $0.2$ and $0.7$, which worked reasonably well for some of our estimators.
@@
Figure 1: Volatility index after Politis et al. (1999) as a function of $m$ for Chatterjee’s correalation (xicor) and Tukey’s shorth for $n=1000$.
@@ 2.2   Choice of $m$
The algorithm has two parameters, the window width $h_{ci}$ for smoothing the interval end points and the window width $h_{\sigma}$ for computing the runnning standard deviation. Politis et al. (1999) recommended $h_{ci}=h_{\sigma}=2$ or $3$, which corresponds to a range of $5$ or $7$ values. To avoid finding a minimum that is too close to $n$, as happens in Fig. 1(b), we have restricted the search range to values $m<n/2$. For efficiency reasons, Politis et al. (1999) recommended not to try every $m$, but only a grid of equidistant values. Our implementation in the function estimate.m.volatility() in the package moonbot therefore only tries out a maximum of 50 values for $m$.
@@
Two other methods have been suggested in the literature, which we have not included in moonboot. Chung and Lee (2001) based their method on the assumptions of root-n consistency and asymptotic normality of the estimator. Although the assumption of root-n consistency might be overcome by replacing factors $n^{1/2}$ with $\tau_{n}$ in the formulas, the normality assumption is essential for that method because it is based on an Edgeworth expansion. In our tests with non-normal estimators, the coverage probabilities were very low and did not even converge to the nominal value for large $n$. Lee and Yang (2020) suggested to use a double bootstrap for estimating the coverage probability for different values of $m$ and to choose the $m$ where this estimate is close to the nominal value. In our experiments, however, the estimated coverage probability based on an in-sample bootstrap was either way too low or almost one and thus rarely close to the nominal value. For xicor, this behaviour was not even monotone in $m$, which made the method difficult to apply. Moreover, it has a runtime complexity of $O(R^{2}n^{2})$, which made Monte Carlo simulations for its evaluation infeasible.
@@ 2.3   Functions provided by moonboot
Like in the boot package, the computation of the confidence interval is split into two steps: the sampling procedure for simulating the bootstrap distribution $T_{m}^{*}$, and the computation of the confidence interval therefrom:
@@
Simulates the bootstrap distribution of the given estimator statistic, which must have been defined as a a function with two arguments: statistic(data, indices). For multidimensonal data, each row in data is assumed to be one data point. Additional arguments are passed to statistic.
@@
Estimates the confidence interval with Eqs. (4) (“basic”) or (5) (“norm”). tau must be a function that computes $\tau_{n}$ from its argument $n$. If it is not provided, it is estimated with estimate.tau() with the default settings of this function. According to the results in Section 4.1, this is not recommended, though, and it is preferrable to provide the correct value for tau.
@@
             beta=seq(0.2, 0.7, length.out=5), method="variance", ...)
Estimates the scaling factor with the method by Bertail et al. (1999). The values for $m$ are tried out as $m_{i}=n^{\beta_{i}}$. min.m is the minimum sample size for which the statistic makes sense and can be computed. The estimation is based on the scaling behaviour of the variance (method="variance") or of the quantile ranges (method="quantile").
@@ 2.4   Typical usage
  boot.stat <- function(indices, dat) {    my.stat(dat[indices])  }
@@
Beware that this automaticlly estimates the asymptotic convergence rate $1/\tau_{n}^{2}$ of the variance of my.stat with the method by Bertail et al. (1999). It is thus better, to provide this rate in the parameter tau, if you happen to know it. For a root-n consistent estimator, e.g., it is $\tau_{n}=\sqrt{n}$ and you can call mboot.ci() as follows [^3] It is more efficient to just set tau=sqrt, of course, but the example uses a more complicated way for the sake of clarity.:
  ci <- mboot.ci(boot.out, tau=function(n) { n^0.5 }, type="basic")
@@ 3   Examples
max(x) $x_{i}\sim\operatorname{unif}(0,1)$ $1$ $n$
moonboot::shorth(x) $x_{i}\sim\operatorname{norm}(0,1)$ 00 $n^{1/3}$
XICOR::xicor(x,y) $x_{i}\sim\operatorname{unif}(-1,1)$ and $y_{i}\sim x_{i}+\mathcal{N}(0,0.5)$ $0.3818147$ $\sqrt{n}$
@@
Table 1: Tested estimators and data generation pocesses. $x=(x_{1},\ldots,x_{n})$ and $y=(y_{1},\ldots,y_{n})$ are the simulated data sets.
@@
The mean is a very well behaved statistic: it is asymptotic normal, root-n consistent and bootstrap consistent. It can thus serve as a simple test case how well the m-out-of-n bootstrap performs for bootstrap consistent estimators. We have simulated data accordng to the density $f(x)=3x^{2}$ for $x\in[0,1]$, which was already used in a study by Dalitz (2017), and for which we have implemented the random number generator rpower() in moonboot.

### v3

@@ 1  Introduction
The bootstrap is a resampling method introduced by Efron (1979), which repeatedly simulates samples of the same size $n$ as the observed data by $n$-fold drawing with replacement. Let us call this method n-out-of-n bootstrap, in order to distinguish it from different sampling schemes. For each of these samples, the estimator $T$ of interest is computed, thereby simulating a sample $T_{1}^{*},\ldots,T_{R}^{*}$ for the distribution of $T$, where $R$ is the number of bootstrap repetitions. For sufficiently smooth estimators $T$, this bootstrap distribution asymptotically approaches the true distribution of $T$ (Giné and Zinn, 1990; Shao and Tu, 1995), and it can thus be used to construct confidence intervals for $T$. The coverage probability of confidence intervals based on flipping the quantiles of the bootstrapped distribution $T_{1}^{*},\ldots,T_{R}^{*}$ at the point estimate (“basic bootstrap”) can be shown to be first order accurate, i.e. up to $O(n^{-1/2})$. Under certain conditions, it is even possible to make the confidence intervals second order accurate, i.e. up to $O(n^{-1})$, by studentized sampling (“bootstrap-tt”) or accelerated bias correction ($BC_{a}$ bootstrap”) (Hall, 1988; DiCiccio and Efron, 1996; Davison and Hinkley, 1997). Compared to other methods for estimating confidence intervals, the bootstrap is more versatile because it neither relies on a likelihood function, nor does it require asymptotic normality. The bootstrap has thus become a standard technique for estimating confidence intervals, and it is even included in vanilla R with the package boot (Canty and Ripley, 2021), which implements five different methods in boot.ci().
@@
There are, however, bootstrap inconsistent estimators, i.e., estimators for which the n-out-of-n bootstrap fails to yield confidence intervals with an asymptotically correct coverage probability. Examples include extreme order statistics (Bickel et al., 1997), the Grenander estimator of a monotonous density (Sen et al., 2010), Chernoff’s estimator of the mode and Tukey’s shorth (Léger and MacGibbon, 2006), or Chatterjee’s rank correlation index (Lin and Han, 2024; Dalitz et al., 2024). See the discussion by Lin and Han (2024) for a literature review and further examples. A possible workaround in these cases is the m-out-of-n bootstrap, which samples only $m<n$ observations. This can be done with or without replacement. Politis and Romano (1994), who introduced this method, only considered sampling without replacement, which they called subsampling. They have shown that the m-out-of-n bootstrap without replacement works under much weaker conditions than the n-out-of-n bootstrap, provided $m$ is chosen such that, for $n\to\infty$, $m\to\infty$ and $m/n\to 0$. The only condition is that the estimator, suitably scaled by some factor $\tau_{n}$, possesses a limit distribution, whereas no smoothness of the estimator or uniformity of the convergence is required. If the sampling is instead done with replacement, Bickel et al. (1997) have shown that a similar result requires additional restrictions on the estimator. Sampling without replacement thus works under weaker conditions than sampling with replacement, and, in the present article, we therefore only consider sampling without replacement. Our package moonboot [^1] This article used version 2.0.x, available, e.g., from https://github.com/cdalitz/moonboot. supports sampling both with and without replacement by means of an argument replace, which is set to FALSE by default.
@@
The wider applicability of the m-out-of-n bootstrap comes at a price, though. One drawback is that the scaling factor $\tau_{n}$ must be known. Bertail et al. (1999) explained that it must be chosen such that $\tau_{n}^{2}\operatorname{Var}(T)$ converges to some constant, and they also suggested a method to estimate the scaling factor with another bootstrap. For root-$n$ consistent estimators, it is $\tau_{n}=\sqrt{n}$, but not all estimators are root-$n$ consistent, especially if they are bootstrap inconsistent, and even if root-$n$ consistency is conjectured it can be difficult to prove (Lin and Han, 2024). Moreover, the m-out-of-n bootstrap can have less than first order accuracy: Politis and Romano (1994) gave a theoretical example with accuracy $O(n^{-1/3})$ for an optimal choice of $m$, and even worse for other choices of $m$. That the accuracy depends on the choice of $m$ was demonstrated, too, in a simulation study by Kleiner et al. (2014). And for the sample quantiles and sampling with replacement, Arcones (2003) has shown that the choice $m\propto n^{2/3}$ is optimal. This raises the problem how to choose $m$, and several heuristics have been suggested in the literature (Politis et al., 1999; Götze and Račkauskas, 2001; Bickel and Sakov, 2008; Chung and Lee, 2001; Lee and Yang, 2020).
@@
Due to these shortcomings, the usual n-out-of-n bootstrap is generally preferable for bootstrap consistent estimators. However, there are bootstrap inconsistent estimators and there is thus need for an R package that facilitates the application of the m-out-of-n bootstrap by providing the required functions. To this end, we present a new package moonboot, evaluate the algorithms implemented therein by means of Monte Carlo simulations, and give recommendations and use cases for their application. The name of the package was inspired by an article by Götze and Račkauskas (2001), who abbreviated the m-out-of-n bootstrap as “moon bootstrap”.
@@ 2  The m-out-of-n bootstrap
Let $X_{1},\ldots,X_{n}$ be i.i.d. random variables, $\theta$ be some parameter of their underlying distribution, and $T_{n}=T_{n}(X_{1},\ldots,X_{n})$ be an estimator for $\theta$. Let us additionally assume [^2] Politis and Romano (1994) did not make this assumption, but it is necessary for the estimation of the scaling factor $\tau_{n}$ with the method by Bertail et al. (1999). that $E(T_{n}^{2})<\infty$. The m-out-of-n bootstrap requires that, for some scaling factor $\tau_{n}$ which behaves for $n\to\infty$ as
@@
Note that condition (1) implies that $\tau_{n}\to\infty$, and thus the convergence (2) requires that $T_{n}$ converges to $\theta$ in probability (in other words: $T_{n}$ must be consistent), because it follows from (2) that
$P(|T_{n}-\theta|>\varepsilon)=F_{n}(-\tau_{n}\varepsilon)+(1-F_{n}(\tau_{n}\varepsilon))\;\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}\;0$
@@
Sherman and Carlstein (2004) observed that the scaling with $\tau_{m}/\tau_{n}$ in (4) can be omitted if the interval is centered around one of the $T_{m}$ instead of $T_{n}$. They suggested to use the first $m$ samples for computing $T_{m}$, but this is an arbitrary choice and thus even the location of the confidence interval is not uniquely determined by the data. Moreover, the lack of scaling increases the interval length. The greater length is compensated by the higher volatility of the location so that the coverage probability still is close to the nominal value. This does not hold, however, for drawing without replacement as $m$ approaches $n$, and a compromise must be made between small interval length and approximate coverage probability. Sherman and Carlstein (2004) suggested a heuristic method for choosing $m$ on basis of a double bootstrap, and they gave the rule of thumb to choose $m$ between $n^{1/2}$ and $n^{2/3}$. However, due to the greater interval length, this method should only be used as a last resort if the convergence rate $\tau_{n}$ is unknown and cannot be estimated from the data. In moonboot, this method is implemented by mboot.ci(..., method="sherman") and the method for estimating $m$ by estimate.m.sherman().
@@ 2.1  Estimation of $\tau_{n}$
and $\beta$ can be estimated with a least-squares fit. For the choice of the test values $m$, Bertail et al. (1999) suggested $m_{i}=n^{\gamma_{i}}$, but wrote that “the difficult problem of choosing the $\gamma_{i}$’s requires more work on a case-by-case basis” [^3] Note that Bertail et al. (1999) wrote $n^{\beta_{i}}$ for the test values $m_{i}$, but we have renamed them to $n^{\gamma_{i}}$ to avoid confusion with the power $\beta$ that is to be estimated.. This casts doubt on the whole method, because removing the necessity of a detailed asymptotic analysis of the estimator is the whole point of estimating $\tau_{n}$ from the data. Indeed, we have not found a sequence of values for the $\gamma_{i}$ that worked equally well in all of our examples. We thus leave the choice of the $\gamma$-sequence as an option to the user and use the default of five values between $0.2$ and $0.7$, which worked reasonably well for some of our estimators.
@@
Figure 1: Volatility index after Politis et al. (1999) as a function of $m$ for Chatterjee’s correlation (xicor) and Tukey’s shorth for $n=1000$.
@@ 2.2  Choice of $m$
The algorithm has two parameters, the window width $h_{ci}$ for smoothing the interval end points and the window width $h_{\sigma}$ for computing the running standard deviation. Politis et al. (1999) recommended $h_{ci}=h_{\sigma}=2$ or $3$, which corresponds to a range of $5$ or $7$ values. To avoid finding a minimum that is too close to $n$, as happens in Fig. 1(b), we have restricted the search range to values $m<n/2$. For efficiency reasons, Politis et al. (1999) recommended not to try every $m$, but only a grid of equidistant values. Our implementation in the function estimate.m.volatility() in the package moonboot therefore only tries out a maximum of 50 values for $m$.
@@
Two other methods have been suggested in the literature, which we have not included in moonboot. Chung and Lee (2001) based their method on the assumptions of root-n consistency and asymptotic normality of the estimator. Although the assumption of root-n consistency might be overcome by replacing factors $n^{1/2}$ with $\tau_{n}$ in the formulas, the normality assumption is essential for that method because it is based on an Edgeworth expansion. In our tests with non-normal estimators, the coverage probabilities were very low and did not even converge to the nominal value for large $n$. Lee and Yang (2020) suggested to use a double bootstrap for estimating the coverage probability for different values of $m$ and to choose the $m$ where this estimate is close to the nominal value. In our experiments, however, the estimated coverage probability based on an in-sample bootstrap was either way too low or almost one and thus rarely close to the nominal value. For xicor, this behavior was not even monotone in $m$, which made the method difficult to apply. Moreover, it has a runtime complexity of $O(R^{2}n^{2})$, which made Monte Carlo simulations for its evaluation infeasible.
@@ 2.3  Functions provided by moonboot
Like in the boot package, the computation of the confidence interval is split into two steps: the sampling procedure for simulating the bootstrap distribution $T_{m}^{*}$, and the computation of the confidence interval therefrom. Using the same interface makes it easy for a user to switch between boot and moonboot. Moreover, this interface is very flexible because it allows for statistics defined on complicated data structures for which only the user knows how to do the sub-indexing. The interface consists in the following two functions:
@@
Simulates the bootstrap distribution of the given estimator statistic, which must have been defined as a a function with two arguments: statistic(data, indices). For multidimensional data, each row in data is assumed to be one data point. Additional arguments are passed to statistic.
@@
Estimates the confidence interval with Eqs. (4) (“basic”) or (5) (“norm”). tau must be a function that computes $\tau_{n}$ from its argument $n$. If it is not provided, it is estimated with estimate.tau() with the default settings of this function. According to the results in Section 4.1, this is not recommended, though, and it is preferable to provide the correct value for tau.
@@
             gamma=seq(0.2, 0.7, length.out=5), method="variance", ...)
Estimates the scaling factor with the method by Bertail et al. (1999). The values for $m$ are tried out as $m_{i}=n^{\gamma_{i}}$. min.m is the minimum sample size for which the statistic makes sense and can be computed. The estimation is based on the scaling behavior of the variance (method="variance") or of the quantile ranges (method="quantile").
@@ 2.4  Typical usage
  boot.stat <- function(dat, indices) {    my.stat(dat[indices])  }
@@
Beware that this automatically estimates the asymptotic convergence rate $1/\tau_{n}^{2}$ of the variance of my.stat with the method by Bertail et al. (1999). It is thus better, to provide this rate in the parameter tau, if you happen to know it. For a root-n consistent estimator, e.g., it is $\tau_{n}=\sqrt{n}$ and you can call mboot.ci() as follows [^5] It is more efficient to just set tau=sqrt, of course, but the example uses a more complicated way for the sake of clarity.:
  ci <- mboot.ci(boot.out, tau=function(n) { nˆ0.5 }, type="basic")
@@ 3  Examples
max(x) $x_{i}\sim\operatorname{unif}(0,1)$ 11 $n$
moonboot::shorth(x) $x_{i}\sim\operatorname{norm}(0,1)$ 0 $n^{1/3}$
XICOR::xicor(x,y) $x_{i}\sim\operatorname{unif}(-1,1)$, $y_{i}\sim x_{i}+\mathcal{N}(0,0.5)$ $0.3818147$ $\sqrt{n}$
@@
Table 1: Tested estimators and data generation processes. $x=(x_{1},\ldots,x_{n})$ and $y=(y_{1},\ldots,y_{n})$ are the simulated data sets.
@@
The mean is a very well behaved statistic: it is asymptotic normal, root-n consistent and bootstrap consistent. It can thus serve as a simple test case how well the m-out-of-n bootstrap performs for bootstrap consistent estimators. We have simulated data according to the density $f(x)=3x^{2}$ for $x\in[0,1]$, which was already used in a study by Dalitz (2017), and for which we have implemented the random number generator rpower() in moonboot.